# Multimodal AI predicts clinical outcomes of drug combinations from preclinical data

### v1

@@ 3.3 Evaluation Metrics
Specifically, for each outcome, given predicted scores $\mathbf{s}=(s_{1},s_{2},\dots,s_{N})$ and corresponding binary labels $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$, Fmax is calculated as:
$F_{\max}=\max_{\tau}F(\tau),$
where the F1 score at threshold $\tau$ is defined as
$F(\tau)=\frac{2\,\text{Prec}(\tau)\,\text{Rec}(\tau)}{\text{Prec}(\tau)+\text{Rec}(\tau)},$
with the precision and recall at threshold $\tau$ being
$\quad\text{Prec}(\tau)=\frac{\sum_{i=1}^{N}y_{i}\,\mathbb{I}(s_{i}\geq\tau)}{\sum_{i=1}^{N}\mathbb{I}(s_{i}\geq\tau)},\quad\text{Rec}(\tau)=\frac{\sum_{i=1}^{N}y_{i}\,\mathbb{I}(s_{i}\geq\tau)}{\sum_{i=1}^{N}y_{i}},$
where $\mathbb{I}(\cdot)$ is the indicator function.
@@ 3.4 Baselines
To test the performance of our proposed Madrigal, we compare Madrigal with six baselines across four modalities on two datasets.
@@ 3.5 Modality Ablation Tests
4 Madrigal-LLM Extension
4.1 Augmenting descriptions of drug combination outcomes
We use GPT-4 to get augmented and diverse safety outcome descriptions [152]. The original descriptions from DrugBank are short and formulaic, which makes the model unable to learn more flexible representations of safety outcome descriptions. We design a prompt that engages GPT-4 using the gpt-4-0125-preview API version with a knowledge cutoff of December 2023 to return the augmented version of each safety outcome description as a medical doctor, pharmacologist, and toxicologist. In the prompt, we ask it not to hypothesize anything about the drug types or methods of action and to restrict it to augmenting just the safety outcome description. There are 269 unique descriptions from DrugBank for the 158 outcomes in the DrugBank dataset, as some classes can have more than one description. We generate three different augmentations for each expert for each class, totaling nine augmented descriptions per class.
We use the following prompt for augmenting descriptions:
Your task is to take original descriptions of drug-drug interactions and augment them. Make sure the information you provide is correct and valid, and the extended text is two sentences long. To ensure diversity in these descriptions, respond as each of the following: medical doctors (physician), pharmacologist and toxicologist. Medical doctors: Especially those specializing in fields such as pharmacology, internal medicine, and clinical pharmacology. Hence, they understand the symptoms and organ systems affected due to the drug-drug interaction. Pharmacologists: Scientists who study the effects of drugs on biological systems. They understand the genetic and molecular mechanism behind these interactions. Toxicologists: Scientists typically work with chemicals and other substances to determine if they are toxic or harmful to humans, other living organisms, or the environment. Your job is to generate nine versions of the original description (given below after ‘Original description’) - three for each profession and at the correct level of understanding and expertise. Your responses for the scientists should be diverse in terms of style, length, and language used. Each of the scientist’s responses must be different in terms of the level of vocabulary used, sentence structure, and word choice. None of the nine responses can be the same or have any repeating sentences. Please also make sure you do not make any assumptions about @Drug1 and @Drug2. An example output is given for you after ‘Example’. The description to augment is given after ‘Task’. Example: Description: The risk or severity of bleeding and bruising can be increased when @Drug1 is combined with @Drug2. Output: {”medical_doctor_1”: ”When @Drug1 is combined with @Drug2, there is an elevated risk of experiencing increased bleeding and bruising, a concern particularly for patients with pre-existing bleeding disorders or those on anticoagulant therapy. These patients should be closely monitored for any signs of abnormal bleeding and adjustments to therapy may be necessary.”, ”medical_doctor_2”: ”The concurrent use of @Drug1 with @Drug2 can lead to a heightened risk or severity of bleeding and bruising, necessitating careful consideration in patients undergoing surgical procedures or those with a history of bleeding problems. Its vital to evaluate the benefits and risks of continuing this drug combination in such contexts.”, ”medical_doctor_3”: ”Combining @Drug1 with @Drug2 can significantly increase the likelihood of bleeding and bruising, an issue of particular concern for individuals with liver impairment or those consuming alcohol regularly. These patients may require dose adjustments or more frequent monitoring to mitigate the risks.”, ”pharmacologist_1”: ”The interaction between @Drug1 and @Drug2 can escalate the risk of bleeding and bruising due to their synergistic effects on inhibiting platelet aggregation or affecting coagulation pathways. Understanding the molecular basis of this interaction is crucial for predicting patient responses and managing therapy effectively.”, ”pharmacologist_2”: ”When @Drug1 and @Drug2 are co-administered, they may interfere with each others pharmacokinetic or pharmacodynamic profiles, leading to an increased risk or severity of bleeding and bruising. This interaction suggests a complex interplay at the cellular level, possibly involving the modulation of enzymes or receptors associated with hemostasis.”, ”pharmacologist_3”: ”The combination of @Drug1 and @Drug2 can intensify bleeding and bruising risks, likely through a mechanism involving the alteration of drug metabolism or the exacerbation of anticoagulant effects. Pharmacologists should explore the specific pathways affected to tailor safer therapeutic strategies.”, ”toxicologist_1”: ”The interaction between @Drug1 and @Drug2 may elevate the risk of bleeding and bruising by enhancing the toxic effects on the bodys hemostatic system, potentially leading to hazardous conditions such as uncontrolled bleeding. Toxicologists should assess the cumulative impact of these drugs to ensure patient safety.”, ”toxicologist_2”: ”Co-administration of @Drug1 with @Drug2 can lead to increased toxicity, manifesting as heightened bleeding and bruising, by overburdening the bodys natural mechanisms for blood clotting. This interaction underscores the importance of toxicological evaluation for identifying potential adverse effects on the hemostatic process.”, ”toxicologist_3”: ”When @Drug1 and @Drug2 are used together, their combined toxic effects can compromise the integrity of the vascular and coagulation systems, increasing the risk of bleeding and bruising. This interaction highlights the need for ongoing monitoring and possibly adjusting dosages to mitigate the toxicological risks involved.”} Task: Description: $\mathsf{DESC}$} Output:
4.2 Large language model with Madrigal
We start with the aligned modality-specific encoders (after first and second-stage optimization) of Madrigal. We use the Mistral 7B [126] model to encode text descriptions of drug-drug interactions. Then, a self-attention layer is applied to the two drug-modality-fused encodings and the text encoding, with an MLP decoder to predict the described safety outcome from the concatenation of self-attention outputs. More specifically,
$z_{\mathrm{d}_{1}}=\textsc{Madrigal}(x_{\mathrm{d}_{1}}),$
$z_{\mathrm{d}_{2}}=\textsc{Madrigal}(x_{\mathrm{d}_{2}}),$
$z_{\mathrm{text}}=\textsc{Mistral}(x_{\mathrm{text}}).$
Then,
$z=\left[z_{\mathrm{d}_{1}},z_{\mathrm{d}_{2}},z_{\mathrm{text}}\right],$
$z_{\mathrm{attn}}=\text{Self-Attention}(z),$
$y=\text{MLP}(z_{\mathrm{attn}}).$
We keep the LLM parameters frozen during training while the weights of the modality-specific encoders, the modality fusion module, the self-attention layer, and the MLP decoder are updated. A binary cross entropy loss with a positive weight $w_{p}$ is used to train the model.
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}[w_{p}y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})]$ (1)
Implementation details.
We trained Madrigal-LLM for 20 epochs using an Adam optimizer with a learning rate 2e-5 and a weight decay of 1e-2. We used a training batch size of 3000 and an evaluation batch of 4000. The weight parameter of the Binary Cross Entropy loss was set to 2.4, where the loss from each positive label is weighted 2.4 times more than each negative label.
4.3 Evaluation of Madrigal-LLM
We designed a baseline that consider semantic similarity to benchmark the performance of Madrigal-LLM, since no other baselines exist in this area. Specifically, we note the outcome for each sample in the evaluation set and compute the top 3 similar outcomes from the training set using PubMed-BERT [153] embedding similarity. For each drug in the top-3 labels, we define the likelihood-weighted prior as the abundance of this drug in the subset of the training set positive samples. Then, for each outcome in the evaluation set, each sample (both positive and negative) is scored by the mean of likelihood-weighted prior to the two drugs. This ensures that the baseline considers the abundance of the drug pairs in the evaluation set with the closest safety outcomes in the training set.
@@ 5 Research Applications of Madrigal
Each DrugBank safety outcomes are annotated with one of nine organs, namely “blood” (hematological), “heart” (cardiovascular), “liver” (hepatic), “kidney” (renal), “gastrointestinal”, “endocrine”, “urinary”, “immune”, “lung”, or otherwise “others/general”. Organs with less than five occurrences (“urinary”, “immune”, “lung”) are not considered in all organ-level analyses. 7 out of 158 safety outcomes, including “adverse effects, decrease”, “cardiotoxicity, decrease”, “hypertension, decrease”, “hypoglycemia, decrease”, “hypotension, decrease”, “nephrotoxicity, decrease”, and “therapeutic efficacy, increase” are considered as potentially beneficial safety outcomes and are thus excluded from all safety-oriented analyses.
@@ 5.1 Drug-Induced Effects on Liver, Heart and QT Prolongation
When correlating our model predictions with annotations in each dataset, since the organ where the toxicity is measured differs, we also make sure the outcomes we consider for our model match such organs (Supplementary Table S1). Specifically, for the DILI (liver) dataset, we obtain predictions from our model with all liver-related outcomes (“excretion rate, increase |||| serum level, decrease |||| efficacy, decrease”, “liver damage, increase”, “liver enzyme elevations, increase”, “metabolism, decrease”, and “metabolism, increase”); for the DICT (cardiovascular) dataset, we obtain predictions from our model with all heart-related outcomes (in total, 44 outcomes); and for the DIQTA (QTc prolongation), we obtain predictions from our model with all QTc prolongation-related outcomes (“QTc prolongation, decrease”, “QTc prolongation, hypotension, increase”, “QTc prolongation, increase”, “QTc prolongation, torsade de pointes, cardiotoxicity, increase”). For DILI and DIQTA, we correlate annotations with predictions for each outcome individually; for DICT, due to the large number of outcomes, we correlate annotations with the average of the highest five predictions across 44 heart-related outcomes.
@@ 5.2 Transporter, Carrier, and Enzyme-Mediated Outcomes
We identify safety outcomes in the DrugBank dataset that are potentially transporter-mediated, including “absorption, decrease”, “absorption, decrease |||| serum level, decrease |||| efficacy, decrease”, “absorption, increase |||| serum level, increase |||| adverse effects, increase”, “excretion rate, decrease |||| serum level, increase”, “excretion rate, increase |||| serum level, decrease |||| efficacy, decrease”, “excretion, decrease”, “excretion, increase”, “serum level of the active metabolites, decrease”, “serum level of the active metabolites, decrease |||| efficacy, decrease”, “serum level of the active metabolites, increase”, “serum level, decrease”, “serum level, increase”. Among them, “excretion rate, decrease |||| serum level, increase”, “excretion, decrease”, “serum level of the active metabolites, increase”, “serum level, decrease”, and “serum level, increase” are considered as safety outcomes that are relevant to increase in serum concentration as explored in [57]. We also identified safety outcomes in the DrugBank dataset that are potentially carrier-mediated, including “absorption, decrease”, “absorption, decrease |||| serum level, decrease |||| efficacy, decrease”, “absorption, increase |||| serum level, increase |||| adverse effects, increase”, “bioavailability, decrease”, “bioavailability, increase”, “protein binding, decrease”, “serum level, decrease”, “serum level, increase”. Similarly, potential enzyme-mediated safety outcomes include “bioavailability, decrease”, “bioavailability, increase”, “metabolism, decrease”, “metabolism, increase”, “protein binding, decrease”, “serum level of the active metabolites, decrease”, “serum level of the active metabolites, decrease |||| efficacy, decrease”, “serum level of the active metabolites, increase”.
@@
5.3 Cancer Combination Therapies
We consider PARPi’s annotated with at least one outcome and have at least five samples in the organ. This left us with the following PARPis: olaparib, talazoparib, and rucaparib. However, the combination safety dataset lacks data for other PARPis, so we cannot analyze them precisely.
To validate the prediction of the model with clinical knowledge about the neutropenia of PARPi combinations, the following safety outcomes are considered to be neutropenia-related outcomes: “neutropenia, increase”, “cytopenia, increase”, “myelosuppression, anemia, severe leukopenia, increase”, and “neutropenia, thrombocytopenia, increase”. To obtain a summary statistics of each PARPi’s general safety profile about interactions with other drugs, we combine each PARPi with all non-PARPi cancer drugs we curate and take the average of the highest three normalized ranks for each organ for each drug pair, representing the potential level of organ-specific safety concern of each drug pair. Here, for each outcome, normalized ranks are computed by ranking Madrigal-predicted scores of each drug pair among scores of all possible drug combinations between DrugBank small molecule compounds in our dataset for this outcome, before normalizing to [0,1]. These drug pair scores are taken directly to plot barplot or boxplot, and the median is taken across these scores for each PARPi and each organ to generate a heatmap.
As a control for all PARPi’s, in relevant analyses, we also include the following drug pairs from [79]:
  • FDA-approved drug combinations (decomposed into pairs).
  • FDA-approved drug combinations (decomposed into pairs) where at least one of the two drugs is on our curated cancer drug list.
  • Clinically investigated drug combinations (decomposed into pairs). These two categories are indistinguishable from the current dataset.
Their scores are calculated in a way similar to PARPi combinations.
@@ 5.4 Type 2 Diabetes Comorbidities
We consider the outcomes in Supplementary Table S6 relevant for myocardial infarction and stroke. We then query Madrigal with pairs composed of pioglitazone or rosiglitazone and all other T2D drugs before taking the mean across all pairs of such drugs for each outcome related to myocardial infarction or stroke.
For hyperkalemia analysis, we consider “hyperkalemia, increase”, “hypotension, hyperkalemia, nephrotoxicity, increase”, “renal failure, hyperkalemia, hypertension, increase”, and “renal failure, hypotension, hyperkalemia, increase” as hyperkalemia-related outcomes. We then query the model with pairs composed of each HF drug with all T2D drugs. We take the maximum value across the above four outcomes as the safety score, representing each drug pair’s level of hyperkalemia-specific safety concern. For each MoA group of HF drugs, all drug pairs containing an HF drug within the group are considered when plotting the point plot, using the geometric mean as the estimator.
@@ 5.5 MASH Combination Therapies
5.6 BeatAML
@@
We used principal component analysis (PCA) to reduce the dimension of gene expression data from 22783 to 150, which retains 90.3% of variance. We binary encode somatic mutation data, considering only pathogenic or potentially pathogenic mutations, and filter out those genes with less than three mutations across all patients. We then used multiple correspondence analysis (MCA) to reduce the dimension of the somatic mutation data from 447 to 30, which retains 93.0% of variance. We also keep clinical attributes with less than 10% missingness and impute with either the most frequent or mean values, depending on whether the attribute is categorical or numeric. We exclude technical and administrative attributes and attributes about patient information after specimen collection. After filtering, the following attributes are kept: “gender”, “ageAtDiagnosis”, “priorMalignancyNonMyeloid”, “cumulativeChemo”, “priorMalignancyRadiationTx”, “priorMDS”, “priorMDSMoreThanTwoMths”, “priorMDSMPN”, “priorMDSMPNMoreThanTwoMths”, “priorMPN”, “priorMPNMoreThanTwoMths”, “riskGroup”, “specificDxAtAcquisition”, “ageAtSpecimenAcquisition”, “specimenGroups”, “specimenType”, “FLT3_ITDCall”, “NPM1Call”, “priorTreatmentTypeCount”, “priorTreatmentRegimenCount”, “priorTreatmentStageCount”.
@@
To adapt Madrigal for personalized drug combination synergy prediction, we first use frozen Madrigal encoders trained with the DrugBank combination safety dataset to generate drug embeddings. Then, we adopt a symmetric bilinear decoder to fuse the two drug embeddings. We concatenate the fused output with dimension-reduced gene expression, somatic mutation data, and clinical attributes before feeding into an MLP, which is trained from scratch to predict binary labels of whether or not drugs combinations are synergistic for the patient, defined above in Methods Sec. 1.9. For each Madrigal model (trained with five seeds), we used five seeds to train the bilinear decoder and the MLP for at most 200 epochs, leading to 25 models evaluated for each group. For simplicity, we do not perform hyperparameter tuning. The model’s hyperparameters are the bilinear decoder output dimension $=128$, MLP hidden dimensions $=[256,128]$, MLP dropout $=0.2$, and learning rate $=0.001$. The AdamW optimizer is used for training.
@@
We used AUROC as our primary performance metric. We calculated AUROC in two distinct ways: (1) Patient-centric AUROC: For each patient, we computed the AUROC across all drug combination predictions, then averaged these values over all patients. (2) Drug-centric AUROC: For each drug combination, we computed the AUROC across predictions from different patients, and then averaged these values. This dual approach provides complementary insights into the model’s performance at both the patient level and the drug combination level.
5.7 Patient-derived xenograft encyclopedia

### v2

@@ 3.3 Performance Metrics
Specifically, for each outcome, given predicted scores $\mathbf{s}=(s_{1},s_{2},\dots,s_{N})$ and corresponding binary labels $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$, Fmax is calculated as: F_max = max_τ F(τ), where the F1 score at threshold $\tau$ is defined as F(τ) = 2  Prec(τ)   Rec(τ)Prec(τ) + Rec(τ), with the precision and recall at threshold $\tau$ being  Prec(τ) = ∑i=1Nyi1(siτ)∑i=1N1(siτ),  Rec(τ) = ∑i=1Nyi1(siτ)∑i=1Nyi, where $\mathbbm{1}(\cdot)$ is the indicator function.
@@ 3.4 Baselines
To test the performance of our proposed Madrigal, we compare Madrigal with six baselines across two modalities on two datasets. These models use either late fusion [51, 52, 53]—where each drug molecule is encoded separately and merged—or early fusion, where molecular interactions are modeled from the start [131, 49, 50].
@@ 4 Research Applications of Madrigal
Each DrugBank safety outcomes are annotated with one of nine organs, namely “blood” (hematological), “heart” (cardiovascular), “liver” (hepatic), “kidney” (renal), “gastrointestinal”, “endocrine”, “urinary”, “immune”, “lung”, or otherwise “others/general” (Supplementary Table S1). Organs with less than five occurrences (“urinary”, “immune”, “lung”) are not considered in all organ-level analyses. 7 out of 158 safety outcomes, including “adverse effects, decrease”, “cardiotoxicity, decrease”, “hypertension, decrease”, “hypoglycemia, decrease”, “hypotension, decrease”, “nephrotoxicity, decrease”, and “therapeutic efficacy, increase” are considered as potentially beneficial safety outcomes and are thus excluded from all safety-oriented analyses.
@@ 4.1 Drug-Induced Effects on Liver, Heart and QT Prolongation
When correlating our model predictions with annotations in each dataset, since the organ where the toxicity is measured differs, we also make sure the outcomes we consider for our model match such organs (Supplementary Table S1). Specifically, for the DILI (liver) dataset, we obtain predictions from our model with all liver-related outcomes (“excretion rate, increase $|$ serum level, decrease $|$ efficacy, decrease”, “liver damage, increase”, “liver enzyme elevations, increase”, “metabolism, decrease”, and “metabolism, increase”); for the DICT (cardiovascular) dataset, we obtain predictions from our model with all heart-related outcomes (in total, 44 outcomes); and for the DIQTA (QTc prolongation), we obtain predictions from our model with all QTc prolongation-related outcomes (“QTc prolongation, decrease”, “QTc prolongation, hypotension, increase”, “QTc prolongation, increase”, “QTc prolongation, torsade de pointes, cardiotoxicity, increase”). For DILI and DIQTA, we correlate annotations with predictions for each outcome individually; for DICT, due to the large number of outcomes, we correlate annotations with the average of the highest five predictions across 44 heart-related outcomes.
@@ 4.2 Transporter, Carrier, and Enzyme-Mediated Outcomes
We identify safety outcomes in the DrugBank dataset that are potentially transporter-mediated, including “absorption, decrease”, “absorption, decrease $|$ serum level, decrease $|$ efficacy, decrease”, “absorption, increase $|$ serum level, increase $|$ adverse effects, increase”, “excretion rate, decrease $|$ serum level, increase”, “excretion rate, increase $|$ serum level, decrease $|$ efficacy, decrease”, “excretion, decrease”, “excretion, increase”, “serum level of the active metabolites, decrease”, “serum level of the active metabolites, decrease $|$ efficacy, decrease”, “serum level of the active metabolites, increase”, “serum level, decrease”, “serum level, increase”. Among them, “excretion rate, decrease $|$ serum level, increase”, “excretion, decrease”, “serum level of the active metabolites, increase”, “serum level, decrease”, and “serum level, increase” are considered as safety outcomes that are relevant to increase in serum concentration as explored in [58]. We also identified safety outcomes in the DrugBank dataset that are potentially carrier-mediated, including “absorption, decrease”, “absorption, decrease $|$ serum level, decrease $|$ efficacy, decrease”, “absorption, increase $|$ serum level, increase $|$ adverse effects, increase”, “bioavailability, decrease”, “bioavailability, increase”, “protein binding, decrease”, “serum level, decrease”, “serum level, increase”. Similarly, potential enzyme-mediated safety outcomes include “bioavailability, decrease”, “bioavailability, increase”, “metabolism, decrease”, “metabolism, increase”, “protein binding, decrease”, “serum level of the active metabolites, decrease”, “serum level of the active metabolites, decrease $|$ efficacy, decrease”, “serum level of the active metabolites, increase”.
@@
4.3 Drug Combination Clinical Trials
We derive pairwise safety comparisons from clinical trial AE data and compare them with those derived from Madrigal predictions for the corresponding drug combinations. We restrict this analysis to AEs with $\geq$5 significant arm pairs across the curated trials, namely, alopecia, anemia, hypoglycemia, and neutropenia.
Agreement between Madrigal predictions and AE data for some adverse event $e$ is assessed by whether the safer arm in the trial also receives the lower Madrigal score. More precisely,
$\mathrm{Agreed}_{e}\;=\;\mathbbm{1}\!\Bigl\{\mathbbm{1}\bigl\{\mathrm{Incid}_{e}^{(1)}\geq\mathrm{Incid}_{e}^{(2)}\bigr\}\;=\;\mathbbm{1}\bigl\{\mathrm{Score}_{e}^{(1)}\geq\mathrm{Score}_{e}^{(2)}\bigr\}\Bigr\}.$
where $\mathbbm{1}\{\cdot\}$ denotes indicator function, $\mathrm{Incid}_{e}^{(i)}$ denotes the incidence of adverse event $e$ for arm $i$, and $\mathrm{Score}_{e}^{(i)}$ denotes the Madrigal predicted score for adverse event $e$. Because both DrugBank and TWOSIDES outcomes can be mapped to some of the AEs, we utilized models trained on both datasets and apply Madrigal trained on DrugBank first to compare predicted scores. If the score difference is fewer than 0.1, we resort to Madrigal trained on TWOSIDES for decision.
@@ 4.4 Type 2 Diabetes Comorbidities
To support our analysis, the T2D medications are sourced from DrugCentral [132] through PrimeKG knowledge graph [108] (from the Orange Book of the US FDA [133]), and supplemented with mechanism of action data from UCSF [114], Mayo Clinic [115], Cleveland Clinic [116], and DrugBank [47] (Supplementary Table S3). We ensure that all medications, if not approved in the US, are marketed in Europe or Japan. Due to our focus on small molecule drugs, insulin and its analogs are excluded from this analysis.
We first query Madrigal with pairs composed of pioglitazone or rosiglitazone and all other T2D drugs before taking the mean across all pairs of such drugs for each outcome related to myocardial infarction or stroke (Supplementary Table S6).
For the hyperkalemia analysis, we consider “hyperkalemia, increase”, “hypotension, hyperkalemia, nephrotoxicity, increase”, “renal failure, hyperkalemia, hypertension, increase”, and “renal failure, hypotension, hyperkalemia, increase” as hyperkalemia-related outcomes. We then query the model with pairs composed of each HF drug with all T2D drugs. We take the maximum value across the above four outcomes as the safety score, representing each drug pair’s level of hyperkalemia-specific safety concern. For each MoA group of HF drugs, all drug pairs containing an HF drug within the group are considered when plotting the point plot, using the geometric mean as the estimator.
@@ 4.5 MASH Combination Therapies
4.6 Drug Combination Synergy Prediction in BeatAML
@@
We use principal component analysis (PCA) to reduce the dimension of gene expression data from 22783 to 150, which retains 90.3% of variance. We binary encode somatic mutation data, considering only pathogenic or potentially pathogenic mutations, and filter out those genes with less than three mutations across all patients. We then use multiple correspondence analysis (MCA) to reduce the dimension of the somatic mutation data from 447 to 30, which retains 93.0% of variance. We also keep clinical attributes with less than 10% missingness and impute with either the most frequent or mean values, depending on whether the attribute is categorical or numeric. We exclude technical and administrative attributes and attributes about patient information after specimen collection. After filtering, the following attributes are kept: “gender”, “ageAtDiagnosis”, “priorMalignancyNonMyeloid”, “cumulativeChemo”, “priorMalignancyRadiationTx”, “priorMDS”, “priorMDSMoreThanTwoMths”, “priorMDSMPN”, “priorMDSMPNMoreThanTwoMths”, “priorMPN”, “priorMPNMoreThanTwoMths”, “riskGroup”, “specificDxAtAcquisition”, “ageAtSpecimenAcquisition”, “specimenGroups”, “specimenType”, “FLT3_ITDCall”, “NPM1Call”, “priorTreatmentTypeCount”, “priorTreatmentRegimenCount”, “priorTreatmentStageCount”.
@@
To adapt Madrigal for personalized drug combination synergy prediction, we first use frozen Madrigal encoders trained with the DrugBank combination safety dataset to generate drug embeddings. Then, we adopt a symmetric bilinear decoder to fuse the two drug embeddings. We concatenate the fused output with dimension-reduced gene expression, somatic mutation data, and clinical attributes before feeding into an MLP, which is trained from scratch to predict binary labels of whether or not drugs combinations are synergistic for the patient, defined above in Methods Sec. 1.9. For each Madrigal model (trained with five seeds), we use five seeds to train the bilinear decoder and the MLP for at most 200 epochs, leading to 25 models evaluated for each group. The model’s hyperparameters are the bilinear decoder output dimension =128=128, MLP hidden dimensions =[256,128]=[256,128], MLP dropout =0.2=0.2, and learning rate =0.001=0.001. The AdamW optimizer is used for training.
@@
For evaluation, 10% of patients or drugs are sampled and all associated responsed data are held out. We use AUROC as our primary performance metric. We calculate AUROC in two distinct ways: (1) Patient-centric AUROC: For each patient, we compute the AUROC across all drug combination predictions, then averaged these values over all patients. (2) Drug-centric AUROC: For each drug combination, we compute the AUROC across predictions from different patients, and then averaged these values. This dual approach provides complementary insights into the model’s performance at both the patient level and the drug combination level.
4.7 Drug Combination Response Prediction in Patient-derived Xenografts